OpenStudy (anonymous):

laplace transform of (a+bt)^2 where a and b are constants

OpenStudy (anonymous):

@oldrin.bataku i see u typing

OpenStudy (anonymous):

same

OpenStudy (anonymous):

tbh dont listen to oldrin he's gonna give u wrong answer and u dont want wrong answer considering ur chinese u should even b on here arent u suppose to be smart

OpenStudy (anonymous):

ok i just need some example.

OpenStudy (anonymous):

$$\mathcal{L}\{(a+bt)^2\}=\mathcal{L}\{a^2+2abt+b^2t^2\}=\frac{a^2}s+\frac{2ab}{s^2}+\frac{2b^2}{s^3}$$

OpenStudy (anonymous):

THT LUK RIGHT

OpenStudy (anonymous):

ya thats right epicChoi i just made sure for you you're all good

OpenStudy (anonymous):

thanks guys i have another question.

OpenStudy (anonymous):

no problem :-) we will help you